A training field is formed by joining a rectangle and two semicircles, as shown below. The rectangle is 97m long and 67m wide. What is the length of a training track running around the field?
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(53937)

.
A training field is formed by joining a rectangle and two semicircles, as shown below.
The rectangle is 97m long and 67m wide. What is the length of a training track running around the field?
~~~~~~~~~~~~~~~~~~

Nothing is shown below.

I really do not understand, for what reason do you pronounce/print these words,
if you do not keep your promising.


I simply suspect that you do not read what you copy-paste-submit.



Answer by math_tutor2020(3838)


It depends if the diagram is this


OR if it's this



If it's the first diagram, then the semicircular portions can be joined together to form a circle of diameter 67 meters.
The circumference would be C = pi*d = 3.14*67 = 210.38 meters approximately.
This would be the distance along the curved portion of the track.
The straight line portions (two copies of 97) are added to get 210.38+97+97 = 404.38

So if your diagram is the first one I've shown above, then the answer is roughly 404.38 meters.
I used the approximation pi = 3.14; use more decimal digits of pi to get a more accurate answer.

If on the other hand the diagram was the second one I've shown above, then you'll have the two semicircular portions combine to a circle of diameter 97 meters.
C = pi*d = 3.14*97 = 304.58 is the curved portion of the track.
304.58+67+67 = 438.58 is the total perimeter.


----------------------------------------------------------------------------------

Summary:
404.38 meters is the approximate perimeter for the first diagram
438.58 meters is the approximate perimeter for the second diagram
Both approximations are based on pi = 3.14